OpenStudy (anonymous):

someone please help, ive been on this problem for like an hour! the length of a rectangle piece is 8 cm more then four time its width. the perimeter is 66 cm. find the length and width

OpenStudy (anonymous):

So if the width is x then the length is 8 + 4x

OpenStudy (anonymous):

perimeter of a triangle: P is 2(width) + 2(length) so we have 2x + 2(8 + 4x) = 66 2x + 16 + 8x = 66 10x = 50 x = 5

OpenStudy (anonymous):

So the width is 5 and the length is : 8 + (4)(5) = 28

myininaya (myininaya):

L=8+4W 66=2L+2W 66=2(8+4W)+2W 66=16+8W+2W -16 -16 50=10W W=5 L=8+4(5)=8+20=28 im too slow

OpenStudy (anonymous):

You can check this result by using this method to solve for perimeter. P = 2(5) + 2(28) which is 10 + 56 = 66 indeed.